OpenStudy (anonymous):

Help please!!!! Find a model for simple harmonic motion if the position at t = 0 is 2, the amplitude is 2 centimeters, and the period is 6 seconds.

OpenStudy (anonymous):

any ideas @MaJosephine

OpenStudy (anonymous):

any ideas @Noura11

OpenStudy (anonymous):

It can be written like this : \[x=a\cos(\omega t+\varphi)\] And we have to determine a, omega and phi. a is the amplitude so : \[a=2 cm=0.02m\] the periode T is : \[T=6s\] but we have : \[\omega =\frac{2\pi}T\] so : \[\omega=\frac{\pi}{3} rd/s\]

OpenStudy (anonymous):

ok i think im following you so far.

OpenStudy (anonymous):

Now to find phi we have that at t=0 , x= 2 cm So : \[0.02=0.02\cos \varphi\] so : \[\cos\varphi=1\] and then : \[\varphi=0\] So the motion time equation is : \[x=0.02\cos\frac{\pi t}{3}\]

OpenStudy (anonymous):

okwhy .oo2 and not just 2

OpenStudy (anonymous):

nvrmnd i got it thank you

OpenStudy (anonymous):

Because, you have to put the length in meters not in centimeters !